Operations Report – January 22nd

Crew 203 Operations Report 22-JAN-2019

SOL: 10

Name of person filing report: Freddy Castaneda

Non-nominal systems: Nothing new to report.

Notes on non-nominal systems: None

Generator (hours run): 9hr 30min; Turned on last night (21Jan2019) at 17:40; Stopped this morning (22Jan2019) at 3:10; Turned on tonight (22Jan2019) at 16:50

Solar SOC – Turned on (21Jan2019) 86%; Turned off (22Jan2019) 100%; Turned on (22Jan2019) 83%

Diesel Reading – Under 1/2 tank.

Propane Reading – About 19%

Ethanol Free Gasoline – Nothing to report.

Water (auxiliary tank) – Not in use.

Water (static tank) – About 35%; 190 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 49%; 147 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 18:40 to refill tank

Water Meter: 01403746

Toilet tank emptied: No.

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Used for EVA to Pooh´s Corner.

Hours: 69.7

Beginning charge: 100%

Ending charge: 86%

Currently charging: Yes

Opportunity rover used: Used for EVA to Pooh´s Corner.

Hours: 46.2

Beginning charge: 100%

Ending charge: 84%

Currently charging: Yes

Curiosity rover used: Used for EVA to Pooh´s Corner.

Hours: 71.6

Beginning charge: 100%

Ending charge: 95%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly.

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Not used.

CrewCar used and why, where? Off campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Suit #3 still pending for review.

Summary of Hab operations: Nothing to report.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report.

EVA Report – January 22nd

EVA # 6

Author: David Mateus

Purpose of EVA:

Show EVA protocols and part of MDRS to the journalist Sole Moller and his photographer Noel Spirandelli

Start time:
11:30

End time:
15:30

Narrative:

The purpose of this EVA was to show our EVA protocols and the most beautiful views of MDRS, so we decided to go to pooh´s corner with rovers and climb one of the hills. After Noel took a lot of pictures we came back to the hab, We parked the rovers at the hab and Oscar had to change his space suit.

Then we walked to Hab Ridge

Destination: we are going to Pooh´s Corner and after Hab Ridge (walking)

Coordinates: 517600 E, 425000 N and 518800 E ,425850 N

Participants:
David Mateus, Oscar Ojeda, Freddy Castañeda, Hermes Bolívar, Sole Moller and Noel Spirandelli

Road(s) and routes per MDRS Map: To get to Pooh´s Corner we are going to take the Cow Dung Road 0110 to the north to get to that corner and Walking to Hab Ridge .

Mode of travel: Walking and Driving

Vehicles used: Spirit, Curiosity and Opportunity

Greenhab Report – January 22nd

Crew 203 Greenhab Report 22-Jan-2019
Greenhab Officer:
Hermes Bolivar

Environmental Control:
Heating
Cooling w/ ambient air (3 hrs)

40% Shade Cloth on
80% Shade Cloth on

Average Temperatures:
Low: 16.7°C
High: 33.2°C

Hours of Supplemental Light:
5

Daily Water Usage of Crops:
15 gallons

Water in Blue Tank (gallons):
50%, 150 Gallons.

Times of Watering for Crops:

10:30
14:07
16:51
19:12

Changes to crops:

* There are two trays available for culture.

Narrative:

I Harvest baby greens and carrot, after the harvest of baby greens and the microgreeens two days ago, i have two trays for culture in the next days.

Harvest:

* 119 g of carrot.
* 109 g of baby greens.

Support/Supplies Needed:

NONE

Sol Summary – January 22nd

Sol: 10 (I can´t believe how fast time passes)

Summary Title: Greetings Martians, we come in peace

Author’s name: Oscar Ojeda

Mission Status: Nominal

Sol Activity Summary: The day began with the long expected visit of Sole Moller and Noel Spirandelli, Journalist and Photographer respectively, who came to document our life on the red planet. We had breakfast before they arrived, greeted them with some coffee and then we began a tour of the station. We then got ready for a short EVA, that took us to a wonderful location, really close to the hab, the Hab Ridge, (So, if you are reading this, I STRONGLY recommend to go there in one of the EVAs). We spend our time doing some recognition of the zone, telling them about the science that is done, and keeping up with our lives on the station. We had a martian lunch with them, and after finishing what was left of the station to be seen, they returned to Earth, and we returned to our mission. We’ll hopefully have at last the movie night Morpheus hasn’t let us have.

Look Ahead Plan: We’re expecting to review some unexpected results on some experiments, to continue with the science, and to do more EVAs to the amazing places we are surrounded by. Entering the 4th quart of the mission, our minds are starting to return home, just a bit ahead of us.

Anomalies in work: None

Weather: Clear skies all day long, low to moderate intensity winds.

Crew Physical Status: In good health.

EVA: 1, a team of 4 + 2 journalists went to the Hab Ridge and Pooh’s Corner

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, Journalist Report, EVA Request, EVA Report.

Support Requested: None